OpenStudy (anonymous):

Please help! will give a medal!!! What is the value of c, correct to one decimal place? A.4.6 B.3.2 C.2.8 D.1.2

OpenStudy (anonymous):



OpenStudy (anonymous):

is one of those angles known? for example, is A a right angle?

OpenStudy (anonymous):

angle C is 24.63

OpenStudy (anonymous):

really? damn now we have to go with the law of cosines

OpenStudy (anonymous):

actually not too bad in this case \[c^2=a^2+b^2-2ab\cos(C)\] is what you need

OpenStudy (anonymous):

therefore \[c^2=5^2+7^2-2\times 5\times 7\cos(24.63)\] and a calculator will get \(c^2\) take the square root to get \(c\)

OpenStudy (anonymous):

its giving me 10.36873858

OpenStudy (anonymous):

the square root is 3.2